How much should you invest each month in order to have $300,000 if your rate of return is 7% compounded monthly and you want to achieve your goal in 40 years? $ How much interest will you earn? $ How much should you invest each month in order to have $300,000 if you want to achieve your goal in 20 years? $ If you deposit the amount you need to achieve your goal in 20 years, how much will your savings be worth after 10 years?
Answer #1

Part 1: We are given the following information

r 7.00%
n 40
frequency 12
FV $3,00,000.00

We need to solve the following equation to arrive at the required FV

So the PV is $ 18,392.03 ie we need to deposit $18,392.03 rounded off today to have $300000 at the end of 40 years @7% compounded monthly.

Interest earned is FV - PV = 300000-18,392.03 =  $2,81,607.97

Part 2) $ How much should you invest each month in order to have $300,000 if you want to achieve your goal in 20 years? $

We are given the following information

r 7.00%
n 20
frequency 12
FV $3,00,000.00

We need to solve the following equation to arrive at the required FV

So the PV is $ 74,280.61 ie we need to deposit $74,280.61 rounded off today to have $300000 at the end of 20 years @7% compounded monthly.

Interest earned is FV - PV = 300000-74,280.61=  $ 2,25,719.39

Part 3) If you deposit the amount you need to achieve your goal in 20 years, how much will your savings be worth after 10 years?

We are given the following information

PV 74,280.61
r 7.00%
n 10
frequency 12

We need to solve the following equation to arrive at the required FV

So the FV is $149278.88 ie after 10 years the FV will by 149278.88